OpenStudy (anonymous):

The sum of four odd integers is 28,458,968. What is the second u?

OpenStudy (anonymous):

Four consecutive odd integers: [2x-1], [2x+1], [2x+3], [2x+5]. Their sum: (2x-1) + (2x+1) + (2x+3) + (2x+5) = 8x + 8 = 8(x+1) = 28,458,968 (x+1) = 28,458,968/8 = 3,557,371 x = 3,557,370 The numbers are: 7,114,739 7,114,741 <= answer 7,114,743 7,114,745

OpenStudy (anonymous):

Impressive :)